### v1

# DynoSurf: Neural Deformation-based Temporally Consistent Dynamic Surface Reconstruction

Abstract
This paper explores the problem of reconstructing temporally consistent surfaces from a 3D point cloud sequence without correspondence. To address this challenging task, we propose DynoSurf, an unsupervised learning framework integrating a template surface representation with a learnable deformation field. Specifically, we design a coarse-to-fine strategy for learning the template surface based on the deformable tetrahedron representation. Furthermore, we propose a learnable deformation representation based on the learnable control points and blending weights, which can deform the template surface non-rigidly while maintaining the consistency of the local shape. Experimental results demonstrate the significant superiority of DynoSurf over current state-of-the-art approaches, showcasing its potential as a powerful tool for dynamic mesh reconstruction. The code is publicly available at https://github.com/yaoyx689/DynoSurf.
Figure 1: Temporally-consistent dynamic meshes (i.e., vertices are corresponded and connections are identical over time) reconstructed by our DynoSurf from time-varying 3D point cloud sequences (i.e., no temporal correspondence) without using any shape-prior, ground-truth surface, and ground-truth temporal correspondence. The color and texture map are used to illustrate correspondence across reconstructed mesh frames.
1 Introduction
The reconstruction of temporally consistent geometric surfaces from a continuous dynamic 3D point cloud sequence without correspondences presents a valuable yet formidable challenge. These geometric surfaces play a crucial role in diverse domains, including film, gaming, animation, virtual reality (VR), augmented reality (AR), and robotics. They facilitate various applications such as motion object editing, texture transfer, and shape analysis, thereby contributing to the creation of realistic visual effects and immersive experiences.
With the development of the static surface reconstruction for a single object from a point cloud kazhdan2006poisson ; shen2021deep ; peng2021shape ; huang2023neural , a natural idea to model the temporally consistent dynamic surfaces is to select a keyframe from the point cloud sequences and reconstruct it as a template surface through static reconstruction technology. Then the non-rigid deformation for the template surface is estimated by aligning it with other frames. High requirements are placed on the selection of the keyframe point cloud with fewer missing parts or occlusions and a similar shape to other frames. Some optimization-based dynamic reconstruction frameworks based on fusion continuously update the template shape during frame-by-frame non-rigid tracking newcombe2015dynamicfusion ; dou2016fusion4d ; slavcheva2017killingfusion . These approaches are effective and pragmatic, but most of them use depth as input and cannot handle the sparse point clouds without any camera information. Some network-based method niemeyer2019occupancy ; tang2021learning ; lei2022cadex no longer define or pre-establish the template shape, but jointly learn dynamic reconstruction. This way may achieve better results but increases the difficulty of decoupling the deformation field and the template space, which makes deformation have less practical significance. Moreover, these methods use ground truth to supervise training and have poor generalization. To address these problems, we first design an adaptive keyframe selection strategy and learn the template surface representation, then combine it with a deformation field to form a dynamic reconstruction framework. We reconstruct the temporally consistent surface by joint learning template surface and deformation field.
Furthermore, the design of the deformation field is particularly important. Optimization-based methods li2008global ; li2009robust ; yao2020quasi often employ the embedded deformation graph sumner2007embedded for non-rigid deformation, but they depend on the spatial position and mesh quality of the template surface. Some methods zhang2023self ; park2021nerfies ; tretschk2021non directly use a network to predict the deformation for each point, lacking explicit control of the local structure. bozic2021neural proposes a neural deformation graph to reconstruct dynamic objects. However, they use this graph to reconstruct the geometric surface and non-rigid deformation simultaneously. This makes it need a more complex loss function, leading to the learning unstable. Therefore, we propose a deformation field based on explicit learnable control points, which improves the explicit controllability of the deformation field based on implicit representation. We also learn the blending weights between the points on the template surface and the control points. This allows us to adaptively learn the deformation consistency of surface points based on motion information.
In this paper, we aim to reconstruct a temporally consistent surface from continuous dynamic 3D point cloud sequences. We propose a reconstruction framework combining a template surface representation and a learnable deformation field based on the learnable control points and blending weight. Extensive experiments on three benchmark datasets demonstrate that our proposed temporally-consistent surface reconstruction outperforms the state-of-the-art methods to a large extent.
To summarize, our contributions are as follows:
  • We propose a new learning framework for reconstructing temporally-consistent dynamic surfaces from time-varying point cloud sequences that can be trained without requiring any shape prior, ground-truth surface, and ground-truth temporal correspondence information as supervision.
  • We propose a coarse-to-fine learning strategy for unsupervised static surface reconstruction based on the deformable tetrahedron representation.
  • We propose a deformation representation based on learnable control points and blending weights that can better maintain local structure while being independent of object category.
2 Related Work
In this section, we briefly review the works closely related to our method, including 3D shape representation, non-rigid shape deformation, and sequential shape reconstruction.
3D Shape Representation.  Generally, 3D shape representation could be divided into two main categories, explicit and implicit, and these can be transformed into one another. The point cloud and triangle mesh are the two most commonly utilized explicit representations for 3D shapes in various applications qi2017pointnet ; li2018pointcnn ; zeng20173dmatch ; hanocka2019meshcnn ; feng2019meshnet . Implicit representation utilizes an isosurface of a field to depict a surface. The Binary Occupancy Field (BOF) kazhdan2013screened ; mescheder2019occupancy ; peng2020convolutional ; boulch2022poco and Signed Distance Field (SDF) kolluri2008provably ; liu2021deep ; park2019deepsdf are two prevalent implicit fields for representing shapes, both dividing the entire space into two regions, inside and outside the shape. The implicit fields could be approximated by a neural network, such as DeepSDF park2019deepsdf , and SAL atzmon2020sal .
Given the shapes explicitly represented as point clouds or triangle meshes, diverse methods exist for their conversion into implicit representations niemeyer2019occupancy ; kazhdan2013screened ; kolluri2008provably ; liu2021deep . Conversely, when transforming implicit shapes into explicit forms, Marching Cubes (MC) MARCHINGCUBE and Marching Tetrahedrons (MT) nielson2008dual stand out as the two widely used methods. Considering the detailed parts of shapes, DMTet shen2021deep and FlexiCubes shen2023flexible introduce offset for each tetrahedron or cube vertex, allowing their methodologies to effectively manage detailed shape components.
Non-rigid Shape Deformation.  Deforming non-rigid shapes presents a challenge due to the high degree of freedom involved in controlling the shape deformation. To mitigate this, methods li2008global ; li2009robust employ an embedded deformation graph sumner2007embedded with local rigid properties for modeling deformation. During optimization, various geometric regularization terms, such as as-conformal-as-possible deformations wu2019global and as-rigid-as-possible deformations yang2019global , are incorporated to impart meaningful deformation. Recently, QNS yao2020quasi and AMM yao2023fast have developed the optimization methods to accelerate the optimization process.
With the advancement of deep learning in 3D vision, numerous learning-based deformation methods have been introduced. DispVoxNet shimada2019dispvoxnets represents source and target shapes using voxel grids, learning an offset field supervised by ground truth offsets. In contrast, FlowNet3D liu2019flownet3d employs PointNet++ qi2017pointnet++ to predict offsets for each point on the shapes. CorrNet3D zeng2021corrnet3d designs a symmetric pipeline to compute deformations between point clouds. RMA-Net feng2021recurrent introduces a new deformation representation with a point-wise combination of several rigid transformations. NDP li2022non proposes a non-rigid motion representation using a pyramid architecture.
Sequential Shape Reconstruction.  Several 3D parametric models, such as SMPL loper2023smpl and SMAL zuffi20173d , have been developed to represent general shapes within specific categories. These parametric models allow for straightforward adjustments in pose and shape by modifying their respective parameters. IPNet chowdhury2022ipnet integrates learned implicit fields with SMPL to control the poses of reconstructed human models. Following this, NPMs palafox2021npms disentangle shapes and poses into latent spaces, enhancing reconstruction accuracy. Conversely, H4D jiang2022h4d and LoRD jiang2022lord initially regress SMPL model parameters to align with each frame of given point clouds. Subsequently, vertex-wise offsets are applied to the resulting SMPL models, approximating the reconstructed shapes to match the provided point cloud sequence. However, these methods are constrained to handling sequences within specific categories, limiting their broad applicability.
Recent studies newcombe2015dynamicfusion ; niemeyer2019occupancy ; tang2021learning ; bozic2021neural have expanded the scope of reconstructing dynamic shapes by predicting deformation fields between different frames. Approaches like Dynamic Fusion newcombe2015dynamicfusion and NDG bozic2021neural establish a deformation graph to capture the deformation fields. O-Flow niemeyer2019occupancy extends ONet mescheder2019occupancy from 3D to 4D, representing deformation fields with a Neural-ODE chen2018neural , while LPDC tang2021learning employs MLPs to model correspondences in parallel. Both O-Flow and LPDC reconstruct the reference shape based on the initial frame, with subsequent frames having their query points mapped to the chosen reference frame. In contrast, I3DMM yenamandra2021i3dmm learns a canonical reference shape for sequences of heads, while CASPR rempe2020caspr and Garment Nets chi2021garmentnets use ground truth canonical coordinates for network training, which is often impractical. Cadex lei2022cadex employs canonical map factorization to represent deformation without requiring correspondences during training, but it may struggle with sequences featuring substantial motions.
3 Proposed Method
3.1 Problem Statement and Method Overview
Let $\mathcal{P}=\{\mathbf{P}_{1},...,\mathbf{P}_{K}\}$ be a time-varying 3D point cloud sequence capturing the geometric changes or motion of scenes/objects over time, where the correspondence across frames is unknown and $\mathbf{P}_{k}\in\mathbb{R}^{N_{k}\times 3}$ is the $k$-th ($k\in[1,~{}K]$) frame containing $N_{k}$ points. We aim to reconstruct from $\mathcal{P}$ a temporally-consistent mesh sequence =$\mathcal{M}=\{\mathbb{M}_{1},...,\mathbb{M}_{K}\}$, where the $k$-th mesh surface $\mathbb{M}_{k}:=\{\mathbf{V}_{k},\mathcal{E}\}$ with $\mathbf{V}_{k}$ and $\mathcal{E}$ being the vertex set and connection set, respectively, the vertices across meshes are corresponded, and all meshes share the same $\mathcal{E}$.
To address this challenging task, we propose an unsupervised learning framework, named DynoSurf, which requires no ground-truth surface and correspondence information as supervision. As illustrated in Fig. 2, DynoSurf mainly consists of two stages: template surface construction (Sec. 3.2) and deformation-based temporal reconstruction (Sec. 4). Generally, we first select a keyframe from $\mathcal{P}$ and reconstruct its surface based on the deformable tetrahedron representation through a coarse-to-fine learning process. Then, we deform the enhanced template surface to simultaneously align with all frames of $\mathcal{P}$ by learning a control points blending-based deformation field through jointly optimizing the template surface and deformation field. In what follows, we will detail the framework.
Figure 2: Illustration of the proposed DynoSurf, which can reconstruct from time-varying point cloud sequences temporally-consistent dynamic surfaces without requiring any ground-truth surface and temporal correspondence information.
3.2 Template Surface Representation via Deformable Tetrahedron
In this stage, we aim to construct a template surface, i.e., a mesh fitting the selected point cloud from $\mathcal{P}$, called keyframe. As illustrated in Fig. 3, we initially select a keyframe, denoted as $\mathbf{P}_{k^{*}}\in\mathcal{P}$, and construct its convex hull as a triangle mesh, referred to as $\mathbb{C}$. Subsequently, we process $\mathbb{C}$ to generate a tetrahedron mesh, denoted as $\mathbb{T}$, where the signed distance field (SDF) of the surface underlying $\mathbf{P}_{k^{*}}$ is defined and learned through a coarse-to-fine learning process. The SDF enables us to extract a mesh, denoted as $\overline{\mathbb{M}}:=\{\overline{\mathbf{V}},\mathcal{E}\}$ with $\overline{\mathbf{V}}$ being the vertex set.
[…]
Figure 3: Illustration of the pipeline of learning template surface via deformable tetrahedron.
Keyframe selection and tetrahedron mesh construction.  Since we will deform the reconstructed template surface from $\mathbf{P}_{k^{*}}$ for temporal reconstruction in the following deformation stage, intuitively, $\mathbf{P}_{k^{*}}$ should be relatively average to facilitate the deformation process. Thus, we determine $k^{*}=\mathop{\arg\min}_{k\in[1,K]}\sum_{l}\texttt{CD}_{\ell_{2}}(\mathbf{P}_{k},\mathbf{P}_{l})$, where $\texttt{CD}_{\ell_{2}}(\cdot,~{}\cdot)$ computes the $\ell_{2}$-norm-based Chamfer distance between two point clouds. We then adopt QuickHull barber1996quickhull to construct the convex hull $\mathbb{C}$ of $\mathbf{P}_{k^{*}}$, which is further dilated and re-meshed to obtain a uniform triangle mesh $\widehat{\mathbb{C}}$ surrounding $\mathbb{C}$. By tetrahedralizing $\widehat{\mathbb{C}}$, we obtain a tetrahedron mesh $\mathbb{T}=\{\mathbf{Q},\mathbf{T}\}$ with $\mathbf{Q}=\{\mathbf{q}_{1},...,\mathbf{q}_{|\mathbf{Q}|}\}$ and $\mathbf{T}=\{\mathbf{T}_{1},...,\mathbf{T}_{|\mathbf{T}|}\}$ being the vertex and tetrahedron sets, respectively. Note that the vertices of $\mathbb{T}$ are distributed both inside and outside of $\mathbb{C}$ due to the dilation and remeshing operations.
Coarse-to-fine learning of deformable tetrahedron-based template surface.  We learn the deformation tetrahedron representation shen2021deep to construct a template surface $\overline{\mathbb{M}}$ that approximates the point surface $\mathbf{P}_{k^{*}}$. Specifically, for each vertex $\mathbf{q}\in\mathbf{Q}$, we train an MLP $F_{\bf{\Phi}}(\cdot)$ parameterized with $\bf{\Phi}$ to learn the SDF value $s(\mathbf{q})$ and the displacement vector $\delta(\mathbf{q})$ with respect to the surface $\overline{\mathbb{M}}$:
$[s(\mathbf{q}),\delta(\mathbf{q})]=F_{\bf{\Phi}}(\Gamma(\mathbf{q})),$ (1)
where $\Gamma(\cdot)$ is the Position Encoding operator. Then, based on the deformed tetrahedral grid points $\{\mathbf{q}+\delta(\mathbf{q})\}_{\mathbf{q}\in\mathbf{Q}}$ and $\{s(\mathbf{q})\}_{\mathbf{q}\in\mathbf{Q}}$, we can employ the differentiable marching tetrahedron algorithm shen2021deep , denoted as $\texttt{DMT}(\cdot,~{}\cdot)$, to extract $\overline{\mathbb{M}}$, i.e., $\overline{\mathbb{M}}=\texttt{DMT}(\{\mathbf{q}+\delta(\mathbf{q})\},\{s(\mathbf{q})\})$.
To train $F_{\bf{\Phi}}(\cdot)$ to encode the SDF of the surface $\overline{\mathbb{M}}$, we adopt the following coarse-to-fine strategy to make the learning process easier and faster: (1) We initialize $F_{\bf{\Phi}}(\cdot)$ by learning the SDF of $\mathbb{C}$, i.e.,
$\min_{\bf{\Phi}}L_{\textrm{init}1}=\sum_{\mathbf{q}\in\mathbf{Q}}\|s(\mathbf{q})-\texttt{SDF}(\mathbf{q},\mathbb{C})\|^{2},$ (2)
where $s(\mathbf{q})$ is the predicted SDF value of $F_{\bf{\Phi}}(\cdot)$ and $\texttt{SDF}(\mathbf{q},\mathbb{C})$ denotes the ground-truth SDF value of convex hull surface $\mathbb{C}$ at $\mathbf{q}$. (2) We then refine $F_{\bf{\Phi}}(\cdot)$ by constraining $\overline{\mathbb{M}}$ and $\mathbf{P}_{k^{*}}$ to be as close as possible, i.e.,
$\min_{\bf{\Phi}}L_{\textrm{init}2}=\tilde{w}_{1}L_{\textrm{CD}}+\tilde{w}_{2}L_{\textrm{Norm}}^{*}+\tilde{w}_{3}L_{\textrm{SDF}},$ (3)
where $L_{\textrm{CD}}:=\texttt{CD}_{\ell_{1}}(\overline{\mathbb{M}},\mathbf{P}_{k^{*}})$ denotes the $\ell_{1}$-norm-based Chamfer distance between $\overline{\mathbb{M}}$ and $\mathbf{P}_{k^{*}}$; $L_{\textrm{Norm}}^{*}:=\texttt{NC}_{\ell_{1}}(\overline{\mathbb{M}},\mathbf{P}_{k^{*}})$ measures the consistency of the corresponding normals (see the Supplementary Material for the definition of $L_{\textrm{Norm}}^{*}$); $L_{\textrm{SDF}}$ is employed to promote the accuracy of the learned SDF, defined as
$L_{\textrm{SDF}}=\frac{1}{|\mathbf{Q}|}\sum_{\mathbf{q}\in\mathbf{Q}}|s(\mathbf{q})-\texttt{SDF}_{{\rm IMLS}}(\mathbf{q},\mathbf{P}_{k^{*}})|,$ (4)
where $\texttt{SDF}_{{\rm IMLS}}(\mathbf{q},\mathbf{P}_{k^{*}})$ is the SDF value at $\mathbf{q}$ approximated through implicit moving least-squares kolluri2008provably :
$\texttt{SDF}_{{\rm IMLS}}(\mathbf{q},\mathbf{P}_{k^{*}})=\frac{\sum_{\mathbf{p}_{j}\in\mathcal{N}(\mathbf{q})}\theta(\|\mathbf{q}-\mathbf{p}_{j}\|,{\zeta})\cdot\langle\mathbf{q}-\mathbf{p}_{j},\mathbf{n}_{j}\rangle}{\sum_{\mathbf{p}_{j}\in\mathcal{N}(\mathbf{q})}\theta(\|\mathbf{q}-\mathbf{p}_{j}\|,\zeta)},$ (5)
where $\mathcal{N}(\mathbf{q})$ denotes the set of nearest points in $\mathbf{P}_{k^{*}}$. $\theta(d,\zeta)=\exp(-d^{2}/\zeta^{2})$ is the weight function. We set $|\mathcal{N}(\mathbf{q})|=10$ and $\zeta=0.1$ by default.
Remark.  It is worth noting that the resulting template surface $\overline{\mathbb{M}}$ will be adaptively updated/enhanced during the following deformation stage to facilitate the deformation stage, thus promoting the quality of reconstructed surfaces of the remaining point cloud frames of $\mathcal{P}$.
3.3 Temporal Reconstruction via Learnable Deformation Field
In this stage, as illustrated in Fig. 4, we propose a learnable deformation field based on control points blending. This deformation field enables us to deform the adaptively-enhanced template surface to achieve temporal reconstruction. By employing this approach, we ensure the temporal consistency of the reconstructed surface sequence.
Figure 4: Illustration of the proposed control points blending-based learnable deformation stage for temporal reconstruction. Note that the adaptively enhanced template surface will be deformed to all frames.
Control points blending-based learnable deformation field.  We first build a set of control point $\mathbf{U}:=\{\mathbf{u}_{1},...,\mathbf{u}_{|\mathbf{U}|}\}$, which are obtained using the farthest point sampling (FPS) algorithm eldar1997farthest on the vertex set of the tetrahedron mesh $\mathbb{T}$ (i.e., $\mathbf{Q}$) such that the control points $\mathbf{u}_{r}$, $r=1,2,\dots,|\mathbf{U}|$ are uniformly distributed in $\mathbb{T}$. Note that the positions of initial control points are also adaptively optimized during training to make the deformation more flexible. We associate each $\mathbf{u}_{r}$ with a rotation matrix $\mathbf{R}_{r}^{k}\in\texttt{SO}(3)$ and a translation $\mathbf{t}_{r}^{k}\in\mathbb{R}^{3}$ for each frame $\mathbf{P}_{k}$, $k=1,2,\dots,K$. Then for each vertex of the template surface $\mathbf{\overline{v}}\in\overline{\mathbf{V}}$, which is the vertex set of the surface $\overline{\mathbb{M}}$. the deformed vertex at the $k$-th frame is represented as
$\mathbf{v}^{k}=\sum_{r=1}^{|\mathbf{U}|}\omega(\mathbf{\overline{v}},\mathbf{u}_{r})\cdot(\mathbf{R}_{r}^{k}\mathbf{\overline{v}}+\mathbf{t}_{r}^{k}),$ (6)
with the weight $\omega(\mathbf{\overline{v}},\mathbf{u}_{r})$ learned via
$\omega(\mathbf{\overline{v}},\mathbf{u}_{r})=\frac{\exp(F_{\bf{\Omega}}([\mathbf{\overline{v}}||(\mathbf{\overline{v}}-\mathbf{u}_{r})]))}{\sum_{r=1}^{|\mathbf{U}|}\exp(F_{\bf{\Omega}}([\mathbf{\overline{v}}||(\mathbf{\overline{v}}-\mathbf{u}_{r})]))},$ (7)
where $\mathbf{F}_{\bf{\Omega}}(\cdot)$ is an MLP parameterized with $\bf{\Omega}$, and [||][$\cdot$|$\cdot$][ ⋅ | | ⋅ ] is the concatenation operator. We initialize $\omega(\mathbf{\overline{v}},\mathbf{u}_{r})$ as
$\tilde{\omega}(\mathbf{\overline{v}},\mathbf{u}_{r})=\frac{\exp(-\|\mathbf{\overline{v}}-\mathbf{u}_{r}\|^{2}/2\eta^{2})}{\sum_{r=1}^{|\mathbf{U}|}\exp(-\|\mathbf{\overline{v}}-\mathbf{u}_{r}\|^{2}/2\eta^{2})},$ (8)
where $\eta=0.1$ by default. The reconstructed surface of $\mathbf{P}_{k}$ can be obtained as $\mathbb{M}_{k}=\{\mathbf{V}_{k},\mathcal{E}\}$ with $\mathbf{V}_{k}$ being the collection of $\mathbf{v}^{k}$.
We also utilize an MLP $F_{\bf{\Xi}}(\cdot)$ parametrized with $\bf{\Xi}$ to learn the deformation transformations of all frames, i.e.,
$[\xi_{r}^{k},\mathbf{t}_{r}^{k}]=F_{\bf{\Xi}}(\Gamma([\mathbf{u}_{r}||k])),$ (9)
where $\xi_{r}^{k}\in\texttt{so}(3)$ is the Lie algebra format of $\mathbf{R}_{r}^{k}$. Owing to the spectral bias property of neural networks rahaman2019spectral , the learned transformations would vary smoothly in both spatial and temporal domains to ensure the spatial and temporal smoothness of the constructed mesh sequence.
Remark. It’s worth noting that our deformation formula is different from the neural deformation graph in bozic2021neural , which only defines the importance weight for each graph node, and blends them according to the spatial positions. Here, we use an MLP to parameterize the weights of any point in space. Additionally, it confines the node positions within the object. These constraints limit the flexibility of the deformation. The surface deformation in RMA-Net feng2021recurrent is defined based on globally shared transformations, which limits the ability to capture the locally correlated characteristics of non-rigid object deformations. By using control points-blending, we can effectively enhance the representation capability of the deformation.
Learning of deformable field with enhanced template surface.  Instead of solely focusing on training the deformation stage, i.e., $F_{\bf{\Omega}}(\cdot)$, $F_{\bf{\Xi}}(\cdot)$ and $\mathbf{U}$, we jointly train the deformation stage and the template surface representation, i.e., $F_{\bf{\Phi}}(\cdot)$ has been warmed-up in the first stage. Specifically, at each iteration, we extract the template surface using $\texttt{DMT}(\cdot,\cdot)$ and then deform it to all frames for temporally consistent reconstruction. Through joint optimization, we aim to enhance the adaptability of the template surface to the deformation process, ultimately leading to improved temporal reconstruction quality. To achieve this goal, we consider the following loss function:
$\min_{\bf{\Phi},\bf{\Omega},\bf{\Xi},\mathbf{U}}L=w_{1}L_{\textrm{R-CD}}+w_{2}L_{\textrm{Norm}}+w_{3}L_{\textrm{R-SDF}}+w_{4}L_{\textrm{Smo}}+w_{5}L_{\textrm{Shape}}.$ (10)
Specifically, $L_{\textrm{R-CD}}$ stands for the proposed robust Chamfer distance to cope with point clouds with holes and outliers, defined as $L_{\textrm{R-CD}}=\frac{1}{K}\sum_{k}L_{\textrm{R-CD}}^{k}$ with
$\footnotesize{L_{\textrm{R-CD}}^{k}=\frac{1}{|\widetilde{\mathbf{V}}_{k}|}\sum_{\mathbf{v}_{i}\in\widetilde{\mathbf{V}}_{k}}\phi_{\alpha}({\mathbf{v}}_{i},\mathbf{p}_{\rho_{i}})\cdot\|{\mathbf{v}}_{i}-\mathbf{p}_{\rho_{i}}\|^{2}+\frac{1}{|\mathbf{P}_{k}|}\sum_{\mathbf{p}_{j}\in\mathbf{P}_{k}}\phi_{\alpha}(\mathbf{p}_{j},{\mathbf{v}}_{\tau_{j}})\cdot\|\mathbf{p}_{j}-{\mathbf{v}}_{\tau_{j}}\|^{2},}$ (11)
where $\widetilde{\mathbf{V}}_{k}$ is the set of points uniformly sampled from $\mathbb{M}_{k}$ with the size of $10^{4}$ by default, and $\phi_{\alpha}(x,y)=\exp(-\alpha\cdot\|x-y\|^{2})$. The normal loss term $L_{\textrm{Norm}}=\sum_{k}\texttt{NC}_{\ell_{1}}(\widetilde{\mathbf{V}}_{k},\mathbf{P}_{k})/K$ constrains the consistency of normals.
The proposed robust SDF loss term $L_{{\textrm{R-SDF}}}$ regularizes the orientation of the deformed mesh, defined as
$L_{\textrm{R-SDF}}=\frac{1}{K\cdot|\mathbf{Q}_{s}|}\sum_{\mathbf{q}\in\mathbf{Q}_{s}}\sum_{k}\overline{\psi}_{\beta}^{k}(\mathbf{q})\cdot|f_{\gamma}(s(\widehat{\mathbf{q}}^{k}))-f_{\gamma}(\texttt{SDF}_{{\rm IMLS}}(\widehat{\mathbf{q}}^{k},\mathbf{P}_{k}))|,$ (12)
where $\widehat{\mathbf{q}}^{k}$ is the deformed positions of $\mathbf{q}$ according to Eq. (6); $\overline{\psi}_{\beta}^{k}(\mathbf{q})=\frac{\psi_{\beta}^{k}(\mathbf{q})}{\sum_{k}\psi_{\beta}^{k}(\mathbf{q})}$ with ${\psi}_{\beta}^{k}(\mathbf{q})=\exp(-\beta\cdot|\texttt{SDF}_{{\rm IMLS}}(\widehat{\mathbf{q}}^{k},\mathbf{P}_{k})|^{2})$ and $\beta=50$ is to measure the confidence of each term. That is, when the deformed point $\widehat{\mathbf{q}}^{k}$ has a smaller value in the SDF approximated by the target point cloud, it is closer to the surface, and we assume that it is more accurate; $\mathbf{Q}_{s}$ is a subset of size $|\mathbf{Q}_{s}|=10^{4}$ randomly sampled from $\mathbf{Q}$. $f_{\gamma}(x)=\frac{1}{1+\exp({-\gamma\cdot x})}$ with $\gamma=10^{2}$ is used to scale the amplified SDF value to $(0,1)$ because we pay more attention to whether the point $\widehat{\mathbf{q}}^{k}$ is inside or outside the surface.
The loss term $L_{\textrm{Smo}}$ ensures the smoothness of the deformation, defined as
$L_{\textrm{Smo}}=\frac{1}{2K|\mathcal{E}|}\sum_{k}\sum_{l\in\mathcal{I}(k)}\sum_{(\mathbf{v}_{i},\mathbf{v}_{j})\in\mathcal{E}}\|({\mathbf{v}}_{i}^{k}-{\mathbf{v}}_{i}^{l})-({\mathbf{v}}_{j}^{k}-{\mathbf{v}}_{j}^{l})\|^{2},$ (13)
where $\mathcal{I}(k)=\{k-1,k+1\}$. When $k=1$ or $k=K$, we only use one neighbor point cloud frame. Finally, it is crucial to ensure that the enhanced template surface maintains a reasonable shape throughout the joint optimization process. Thus, we constrain the deformation field of the template surface as unit transformation as possible with a loss term $L_{\textrm{Shape}}$:
$L_{\textrm{Shape}}=\sum_{\mathbf{u}_{r}\in\mathbf{U}}\|[\xi_{r}^{k^{*}},\mathbf{t}_{r}^{k^{*}}]\|^{2},$ (14)
where $[\xi_{r}^{k^{*}},\mathbf{t}_{r}^{k^{*}}]\in\mathbb{R}^{6}$ denotes the rotation matrix represented by Lie algebra format and translation vector of the keyframe.
4 Experiments
4.1 Experimental Settings
Implementation details.  In our approach, we employed Multi-Layer Perceptrons (MLPs) consisting of 5 linear layers with a feature dimension of 128 to implement $F_{\bf{\Phi}}(\cdot)$, $F_{\bf{\Xi}}(\cdot)$, and $F_{\bf{\Omega}}(\cdot)$. The default weights setting in the loss function are listed in Tab. 1. Training was conducted using the ADAM optimizer kingma2014adam with a learning rate of $1\times 10^{-4}$. We performed all experiments on a single NVIDIA A6000 GPU.
Table 1: Default parameter settings and runtime. For the learning template surface, we listed iterations and running time for coarse/fine stage.
Learning template surface Temporal reconstruction
$\tilde{w}_{1}$ $\tilde{w}_{2}$ $\tilde{w}_{3}$#\##Iters Time $w_{1}$ $w_{2}$ $w_{3}$ $w_{4}$ $w_{5}$#\##Iters Time
500 0.001 50 $10^{3}$/5000 10 s/2 mins 500 0.001 100 $10^{3}$ 1 $10^{4}$ 30 mins
Datasets.  We utilized three well-established benchmark datasets, namely DFA-UST bogo2017dynamic , DT4D li20214dcomplete , and AMA vlasic2008articulated , for our evaluation. From motion sequences, we selected 17 consecutive frames to form a set of input point clouds. Randomly sampling 5000 points from the ground truth mesh, we ensured diversity in our point cloud selection. Adopting the train and test data splits provided by lei2022cadex , we used 109 sets and 89 sets from the test data of the DFAUST dataset and DT4D-animal dataset, respectively. The AMA dataset comprises 10 motion sequences; for supervised learning comparison methods, we employed 1176 sets from 7 sequences as the train set and 32 sets of point clouds from the remaining 3 sequences ("crane", "march1" and "samba") as the test set. To maintain consistency across sequences, we normalized the diagonal length of the bounding box in the first frame to 1 and centered it at the origin of the coordinates. Subsequent point clouds underwent transformations based on the scale and translation of the first frame. We utilized Meshlab’s point cloud estimation normal method for normal estimation and adjusted the normal directions to face outward.
Refer to caption(a) Input(b) LPDC tang2021learning (c) Cadex lei2022cadex (d) Ours00.06
Figure 5: Comparison of visual results by different methods on the DFAUST dataset.
Table 2: Quantitative comparisons of different methods on three datasets.
Dataset Method CD (×$\times 10^{-4}$) $\downarrow$ NC $\uparrow$ F-$0.5\%$uparrow F-$1\%$uparrow Corr. (×$\times 10^{-2}$) $\downarrow$
DFAUST bogo2017dynamic LPDC tang2021learning 2.430 0.929 0.299 0.633 1.46
Cadex lei2022cadex 1.062 0.941 0.519 0.823 1.28
Ours 0.688 0.953 0.894 0.985 1.02
DT4D li20214dcomplete Cadex lei2022cadex 22.377 0.868 0.269 0.532 4.99
Ours 1.036 0.933 0.582 0.870 3.23
AMA vlasic2008articulated LPDC tang2021learning 108.822 0.666 0.045 0.101 14.1
Cadex lei2022cadex 71.041 0.663 0.052 0.119 13.6
Ours 0.320 0.918 0.943 0.683 4.44
Refer to caption(a) Input(b) LPDC tang2021learning (c) Cadex lei2022cadex (d) Ours00.15
Figure 6: Comparison of visual results by different methods on the AMA dataset
Methods under comparison.  We conducted a comparative analysis involving our proposed method and two cutting-edge techniques specialized in dynamic surface reconstruction from sequences of point clouds: LPDC tang2021learning and Cadex lei2022cadex . Note that both Cadex and LPDC necessitate ground-truth occupancy values and ground-truth correspondences for supervision during training. For Cadex, we employed the officially released pre-trained model provided by the authors for testing on DFAUST and DT4D datasets. As for LPDC, the authors exclusively performed experiments on the DFAUST dataset and released the pre-trained model. To ensure a comprehensive comparison, we introduced the AMA dataset, a human body dataset characterized by significantly larger deformation and movement ranges compared to DFAUST. Furthermore, we retrained Cadex and LPDC on the AMA dataset.
Evaluation metrics.  Following previous works mescheder2019occupancy ; lei2022cadex ; peng2021shape , we scaled the test data to its original size and quantitatively assessed various methods using $\ell_{2}$-norm-based Chamfer Distance (CD), Normal Consistency (NC), F-score with thresholds of $0.5\%$ (F-$0.5\%$) and $1\%$ (F-$1\%$), and Correspondences Error (Corr.) to evaluate different methods quantitatively. For the visual results, we assigned the same colors to corresponded vertices to indicate the accuracy of learned correspondences by different methods. In addition, we also visualized the pointwise distance error from the ground-truth surface.
4.2 Comparisons with State-of-the-Art Methods
Results of clean and complete data.  Table 2 shows the comparison of our method against state-of-the-art methods on the three datasets. The results demonstrate that our method attains the highest accuracy. The visual comparisons in Figs. 5 and 6 further validate the significant superiority of our method. Notably, Cadex and LPDC exhibit notably poor performance on the AMA dataset. This discrepancy could be attributed to the smaller data volume of the AMA dataset compared to DFAUST and DT4D, shedding light on the limitations of these supervised learning methods. We also refer readers to the Supplementary Material and Video Demo for more visual results.
Results of noisy and partially missing data.  We also evaluate the performance of our method on noisy data. We add $0.5\%$ Gaussian noise for each point in a point cloud sequence in DFAUST dataset bogo2017dynamic and show the results in Fig. 7. Since point cloud data collected by real-depth cameras often contain missing parts, we tested the situation when there are some holes in the surface. For a sequence in "swing" motion sequences of AMA dataset vlasic2008articulated , we placed two virtual cameras in front of and behind the mesh model for simulated acquisition and merged the point clouds scanned by the two cameras. Fig. 7 shows the performance of our method and the variants. We can see that our method can reconstruct partially missing meshes. More results are presented in the Supplementary Material.
Figure 7: Results by our method on noisy data (top) and partially missing data (bottom).
4.3 Ablation Study
Refer to caption(a)(b)(c)(d)(e)(f)(g)
Figure 8: Comparison of visual results by different variants of the deformation field in our method. (a) Input; (b) Point-wise deformation based on an MLP; (c) Fixed blending weight formula; (d) Fixed control point positions; (e) $|\mathbf{U}|=10$; (f) $|\mathbf{U}|=100$; (g) Ours.
In this section, we conducted thorough ablation studies targeting key components of our framework on the AMA dataset, where we took 1 set from each sequence and 10 sets in total for data diversity. We focused on the following aspects:
Table 3: Ablation studies of our method on the AMA dataset vlasic2008articulated . The best and second best results are highlighted in bold and underlined.
Variants CD(×$\times 10^{-4}$) $\downarrow$ NC $\uparrow$ F-$0.5\%$uparrow F-$1\%$uparrow Corr. (×$\times 10^{-2}$) $\downarrow$
MLP-Deform 0.412 0.909 0.917 0.618 5.91
w. $\tilde{\omega}(\mathbf{\overline{v}},\mathbf{u}_{r})$ 0.475 0.911 0.890 0.564 3.26
Fixed $\{\mathbf{u}_{r}\}$ 0.372 0.911 0.927 0.637 3.23
w.o. $L_{\textrm{R-SDF}}$ 0.330 0.918 0.937 0.649 3.14
w.o. $L_{\textrm{Norm}}$ 0.656 0.900 0.884 0.582 3.30
w.o. $L_{\textrm{Smo}}$ 0.378 0.914 0.923 0.616 3.30
w.o. $L_{\textrm{Shape}}$ 0.335 0.919 0.936 0.636 2.91
$|\mathbf{U}|=10$ 0.437 0.910 0.910 0.594 3.30
$|\mathbf{U}|=30$ (Ours) 0.324 0.919 0.939 0.649 2.89
$|\mathbf{U}|=100$ 0.315 0.919 0.943 0.651 3.16
1) Learnable deformation field.  To evaluate the effectiveness of the proposed deformation field, we conducted experiments by maintaining the other components and substituting the deformation field:
  • Pointwise deformation based on an MLP: For any point $\overline{\mathbf{v}}\in\mathbb{R}^{3}$ in the template surface and frame index $k$, we established a deformation function $F_{\bf{\Psi}}(\Gamma([\overline{\mathbf{v}}||k]))$ and obtained the deformed position $\mathbf{v}^{k}=\overline{\mathbf{v}}+F_{\bf{\Psi}}(\Gamma([\overline{\mathbf{v}}||k]))$. Here, $F_{\bf{\Psi}}(\cdot)$ is an MLP with 5 linear layers and a feature dimension of 128.
  • Fixed blending weight formula: We replaced the learnable blending weights in Eq. (6) by setting it to the value computed by $\tilde{\omega}(\mathbf{\overline{v}},\mathbf{u}_{r})$ in Eq. (8) throughout the learning process.
  • Fixed control point positions: We maintained the positions ${\mathbf{u}_{r}}$ of the control points at their initial values.
From Tab. 3 and Fig. 8, the three variants produce worse quantitative and visual results than Ours, verifying the advantage of our proposed deformation field. Fig. 9 also visualizes the learned weights via Eq. (7) and those computed via Eq. (8) for blending, where we can see that the learned weights can capture the motion information and well handle the deformations for the parts with close spatial locations but inconsistent motions.
Refer to caption(a) Input(b) Computed by Eq. (8)(c) Ours
Figure 9: Visual learnable control points and blending weights. We assigned a color $\mathbf{c}_{r}$ to each control point $\mathbf{u}_{r}$ based on its spatial position, and obtained the color for each $\mathbf{\overline{v}}$ by $\sum_{r=1}^{|\mathbf{U}|}\omega(\mathbf{\overline{v}},\mathbf{u}_{r})\cdot\mathbf{c}_{r}$. Closer colors indicate more similar deformation fields.
2) Size of $|\mathbf{U}|$.  From Tab. 3 and Fig. 8, where we set different numbers of control points i.e., 10, 30, and 100, we can observe that too few control points result in insufficient degrees of freedom and suboptimal alignment, while too many control points can weaken local consistency. We set $|\mathbf{U}|=30$ as a balanced trade-off.
3) Loss functions during temporal reconstruction.  To assess the effectiveness of various loss terms used in the temporal reconstruction stage, we individually excluded the robust SDF loss $L_{\textrm{R-SDF}}$, deformation smooth loss $L_{\textrm{Smo}}$, and shape maintenance loss $L_{\textrm{Shape}}$, while keeping other aspects unchanged. The results presented in Tab. 3 demonstrate their necessity and contributions.
4) Keyframe selection and template surface learning.  Compared to directly setting the middle frame as a keyframe, Fig. 10 shows that the selected keyframe by our strategy is closer to the average shape. After learning the template surface, we also show the results of only using the fine stage (Fig. 10 (b)) and using the coarse-to-fine stage (Fig. 10 (c)). Based on Fig. 10 (c), we further show the enhanced template surface after the temporal reconstruction process.
Refer to caption(a) Input: 1-st5-th9-th14-th16-th17-th(b) w.o. Coarse(c) Coarse-to-fine(d) Enhanced
Figure 10: Visual results of the template surface learned from different settings. For input point clouds, we use the green circle to mark the middle frame and the red circle to mark the template frame we selected.
5 Conclusion and Discussion
We introduced a new unsupervised temporally-consistent dynamic surface reconstruction framework for time-varying point cloud sequences. Technically, we integrated template surface representation based on the deformable tetrahedron and a learnable deformation field. We first proposed a coarse-to-fine learning strategy for constructing the template surface. Then we designed a learnable deformation field using the learnable control points and blending weights for temporally reconstruction. We jointly optimized the enhanced template surface and the learnable deformation field. Experiments demonstrate that our proposed method performs even much better than SOTA supervised learning methods.
Figure 11: Failure case of our method.
Although our method has achieved good performance on many examples, it may encounter limitations in cases where the point cloud in certain areas is sparse or the local structure is elongated but exhibits significant non-rigid deformations (as depicted in the Fig. 11). This is attributed to our utilization of Chamfer distance as the alignment supervision, which restricts the establishment of correspondences between adjacent frames solely based on spatial proximity. Incorporating more robust metrics for shape alignment would enhance the effectiveness of our approach.
Supplementary Material
More visual results
We present more visual results in this section and showcase the complete action sequences in the supplementary video.
Comparisons with state-of-the-art methods.  We present additional visual results for comparisons with LPDC [50] and Cadex [21] on the DFAUST dataset [3] (Fig. 12), AMA dataset [52] (Fig. 13) and DT4D dataset [25] (Fig. 15). It is evident that our method outperforms other approaches.
Refer to caption(a) Input(b) LPDC [50](c) Cadex [21](d) Ours00.06
Figure 12: Comparison of visual results by different methods on the DFAUST dataset.
Refer to caption(a) Input(b) LPDC [50](c) Cadex [21](d) Ours00.15
Figure 13: Comparison of visual results by different methods on the AMA dataset [52].
Refer to caption(a)(b)(c)(d)(e)(f)
Figure 14: Comparison of visual results by different variants of excluding a certain loss in our method. (a) Input; (b) w.o. $L_{\textrm{R-SDF}}$; (c) w.o. $L_{\textrm{Norm}}$; (d) w.o. $L_{\textrm{Smo}}$; (e) w.o. $L_{\textrm{Shape}}$; (f) Ours.
Refer to caption(a) Input(b) Cadex [21](c) Ours00.0600.200.1
Figure 15: Comparison of visual results by different methods on the DT4D dataset.
Ablation study.  Additionally, Fig. 14 illustrates the visual results of our method when different loss terms are omitted. We can see that our method, incorporating all loss terms, achieves the best results.
Performance of our method on noisy data and partially missing data.  Furthermore, in Fig. 16, we showcase more visual results of our method applied to noisy data constructed from the DFAUST dataset [3] and partially missing data constructed from AMA dataset [52], respectively.
Figure 16: Visual results of our method on constructed noisy data from the DFAUST dataset [3] (Top) and partially missing data from AMA dataset [52] (Bottom).
Technical Details
In this section, we explain the definitions for some loss functions and evaluation metrics mentioned in the paper. The Chamfer Distance between $\mathbf{X}$ and $\mathbf{Y}$ is defined as
$\texttt{CD}_{\ell_{p}}(\mathbf{X},\mathbf{Y})=\frac{1}{|\widetilde{\mathbf{X}}|}\sum_{\mathbf{x}\in\widetilde{\mathbf{X}}}\left\|\mathbf{x}-\hat{\mathbf{y}}\right\|_{\ell_{p}}+\frac{1}{|\widetilde{\mathbf{Y}}|}\sum_{\mathbf{y}\in\widetilde{\mathbf{Y}}}\left\|\mathbf{y}-\hat{\mathbf{x}}\right\|_{\ell_{p}},$ (15)
where $p=1,2$. $\hat{\mathbf{y}}\in\widetilde{\mathbf{Y}}$ and $\hat{\mathbf{x}}\in\widetilde{\mathbf{X}}$ are the closest points for the $\mathbf{x}$ and $\mathbf{y}$ respectively. When $\mathbf{X}(\mathbf{Y})$ is a point set, $\widetilde{\mathbf{X}}=\mathbf{X}(\widetilde{\mathbf{Y}}=\mathbf{Y})$. When $\mathbf{X}(\mathbf{Y})$ represents a mesh, $\widetilde{\mathbf{X}}(\widetilde{\mathbf{Y}})$ denotes the sampling point on the mesh. In the learning phase, we set the number of sampling points to $10^{4}$, and during the evaluation, we set the number of sampling points to $10^{5}$.
In Eqs. (3) and (10), the normal consistency of $\mathbf{X}$ and ${\mathbf{Y}}$ is defined as
$\texttt{NC}_{\ell_{1}}(\mathbf{X},\mathbf{Y})=\frac{1}{|\mathbf{N}_{x}|}\sum_{\mathbf{n}_{x}\in\mathbf{N}_{x}}\left|1-|\langle\mathbf{n}_{x},\mathbf{n}_{\hat{y}}\rangle|\right|+\frac{1}{|\mathbf{N}_{y}|}\sum_{\mathbf{n}_{y}\in\mathbf{N}_{y}}\left|1-|\langle\mathbf{n}_{y},{\mathbf{n}}_{\hat{x}}\rangle|\right|,$ (16)
[…]
Following [31, 38, 44], the F-score to evaluate the reconstruction accuracy is defined as:
[…]
where
$\textrm{Recall}(\mathbb{M}_{k},\mathbb{M}_{k}^{\textrm{GT}},\epsilon)=\left|\left\{\mathbf{v}_{1}\in\widetilde{\mathbf{V}}_{k},\textrm{s.t.}\min_{\mathbf{v}_{2}\in\widetilde{\mathbf{V}}_{k}^{\textrm{GT}}}\|\mathbf{v}_{1}-\mathbf{v}_{2}\|<\epsilon\right\}\right|,$
$\textrm{Precision}(\mathbb{M}_{k},\mathbb{M}_{k}^{\textrm{GT}},\epsilon)=\left|\left\{\mathbf{v}_{2}\in\widetilde{\mathbf{V}}_{k}^{\textrm{GT}},\textrm{s.t.}\min_{\mathbf{v}_{1}\in\widetilde{\mathbf{V}}_{k}}\|\mathbf{v}_{1}-\mathbf{v}_{2}\|<\epsilon\right\}\right|.$
[…]
We denote $\{\mathbf{V}^{\textrm{GT}}(k)\}_{k}$ as the temporally consistent corresponding points on the ground-truth surfaces. We set $|\mathbf{V}^{\textrm{GT}}(k)|=10^{5}$. In the first frame, we construct the index set of nearest points:
$\texttt{Index}=\{\rho_{i}|\mathbf{v}_{\rho_{i}}=\mathop{\arg\min}_{\mathbf{v}\in\mathbf{V}_{1}}\|\mathbf{v}-\mathbf{v}_{i}^{\textrm{GT}}(1)\|\},$
where $\mathbf{V}_{1}$ is the vertex set of $1$-st reconstructed mesh $\mathbb{M}_{1}$, $\mathbf{v}_{i}^{\textrm{GT}}(1)\in\mathbf{V}^{\textrm{GT}}(1)$. Following [34, 21], the Correspondences Error is defined as
[…]